OpenStudy (danielbarriosr1):

473.25=142.10+8.95h Solve for h

OpenStudy (anonymous):

Are you there daniel?

OpenStudy (danielbarriosr1):

yes

OpenStudy (anonymous):

Alright so whenever you're solving for a variable, the objective is to isolate it. What can we do to isolate the h in this case?

OpenStudy (anonymous):

isolate, to get it by itself.

OpenStudy (anonymous):

473.25=142.10+8.95h What is getting in the way of isolating h?

OpenStudy (danielbarriosr1):

substract by 142.10 then divide by 8.95 to leave h alone

OpenStudy (anonymous):

Exacto-o. Tell me what you get.

OpenStudy (anonymous):

Are you dividing by hand?

OpenStudy (danielbarriosr1):

i am @Lethal

OpenStudy (danielbarriosr1):

it is the only thing i need. Please

OpenStudy (anonymous):

Use a calculator. I don't see why your teacher wouldn't let you use one for arithmetic.

OpenStudy (danielbarriosr1):

she does

OpenStudy (danielbarriosr1):

but i dont have one here

OpenStudy (anonymous):

473.25=142.10+8.95h -142.10 -142.10 331.15=8.95/h Use a calculator.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Okay. Well you know what to do then, but it's 37

OpenStudy (anonymous):

That's the answer.

OpenStudy (anonymous):

h=37

OpenStudy (danielbarriosr1):

thanks